# Online Combinatorial Optimization with Graphical Dependencies

### v1

@@ 5 Lower Bounds for Minimization Problems
In the hardness instance, our underlying graph is generated inductively for $k\geq 0$ as follows: Define $G_{0}=K_{2}$. For each $i\in[k]$ define $G_{i}$ as the graph which results from replacing each edge $uv\in E(G_{i-1})$ with two disjoint paths of length two, i.e., remove edge $uv$ and add new vertices $x,y$ along with new edges $ux,xv,uy,yv$. Each pair of vertices $x,y$ born from the same edge $uv\in E(G_{i-1})$ in this way are called twins. For each $v\in V(G_{k})$, the rank of $v$ is the first $i$ in which $v\in V(G_{i})$. Moreover, the parent of each vertex $v\in V(G_{k})$ of rank $i\geq 2$ is defined as the unique node $p$ of rank $i-1$ adjacent to $v$ in $G_{i}$. If $p$ is the parent of $v$, we call $v$ the child of $p$. Notice that each vertex $v\in V(G_{k})$ of rank $i\in[k-1]$ has two pairs of twin children.
For our instance of online Steiner tree, our underlying graph is $G_{k}$. We set one vertex of rank 00 as the root $r$, and the other vertex of rank 00 as the first arrival $X_{1}$. The second arrival $X_{2}$ is chosen uniformly from one of the two twin vertices of rank $1$. For the remainder of arrivals, the adversary does the following: Let $\{x,y\}$ be the pair of twin vertices closest to $r$ in $G_{k}$ for which
@@ A.1 Hardness for $p$-Sample Maximization Problems
Now, we return to consider the setting where a $p$-fraction of values $X_{i}$ are revealed up front as a sample. Suppose $k^{\prime}$ is the index of the the largest non-zero sample $X_{k^{\prime}}=M^{k^{\prime}-1}$. It is not hard to see that the optimal online algorithm should wait at least until item $X_{k^{\prime}}$ before stopping. Moreover, since each sample revealed after $X_{k^{\prime}}$ is 00, the conditional distribution of values $X_{k^{\prime}+1},\dots,X_{n}$ given the sample is only worse than it is in the setting of ${\rm ALG}_{k^{\prime}}$ above. Hence, the algorithm cannot obtain expected value better than ${\rm ALG}_{k^{\prime}}$, so ${\mathbb{E}}[{\rm ALG}\mid k^{\prime}]\leq{\rm ALG}_{k^{\prime}}=M^{k^{\prime}-1}$. Therefore, we have

### v2

@@ 5 Lower Bounds for Minimization Problems
In the hardness instance, our underlying graph is generated inductively for $k\geq 0$ as follows: Define $G_{0}=K_{2}$. For each $i\in[k]$ define $G_{i}$ as the graph which results from replacing each edge $uv\in E(G_{i-1})$ with two disjoint paths of length two, i.e., remove edge $uv$ and add new vertices $x,y$ along with new edges $ux,xv,uy,yv$. Each pair of vertices $x,y$ born from the same edge $uv\in E(G_{i-1})$ in this way are called twins. For each $v\in V(G_{k})$, the rank of $v$ is the first $i$ in which $v\in V(G_{i})$. Moreover, the parent of each vertex $v\in V(G_{k})$ of rank $i\geq 2$ is defined as the unique node $p$ of rank i1i-1 adjacent to $v$ in $G_{i}$. If $p$ is the parent of $v$, we call $v$ the child of $p$. Notice that each vertex $v\in V(G_{k})$ of rank $i\in[k-1]$ has two pairs of twin children.
For our instance of online Steiner tree, our underlying graph is $G_{k}$. We set one vertex of rank 0 as the root $r$, and the other vertex of rank 0 as the first arrival $X_{1}$. The second arrival $X_{2}$ is chosen uniformly from one of the two twin vertices of rank $1$. For the remainder of arrivals, the adversary does the following: Let $\{x,y\}$ be the pair of twin vertices closest to $r$ in $G_{k}$ for which
@@ A.1 Hardness for $p$-Sample Maximization Problems
Now, we return to consider the setting where a $p$-fraction of values $X_{i}$ are revealed up front as a sample. Suppose $k^{\prime}$ is the index of the the largest non-zero sample $X_{k^{\prime}}=M^{k^{\prime}-1}$. It is not hard to see that the optimal online algorithm should wait at least until item $X_{k^{\prime}}$ before stopping. Moreover, since each sample revealed after $X_{k^{\prime}}$ is 0, the conditional distribution of values $X_{k^{\prime}+1},\dots,X_{n}$ given the sample is only worse than it is in the setting of ${\rm ALG}_{k^{\prime}}$ above. Hence, the algorithm cannot obtain expected value better than ${\rm ALG}_{k^{\prime}}$, so ${\mathbb{E}}[{\rm ALG}\mid k^{\prime}]\leq{\rm ALG}_{k^{\prime}}=M^{k^{\prime}-1}$. Therefore, we have